# A Semantic Search Engine for Mathlib4

### v1

@@ Abstract
The interactive theorem prover, Lean, enables the verification of formal mathematical proofs and is backed by an expanding community. Central to this ecosystem is its mathematical library, mathlib4, which lays the groundwork for the formalization of an expanding range of mathematical theories. However, searching for theorems in mathlib4 can be challenging. To successfully search in mathlib4, users often need to be familiar with its naming conventions or documentation strings. Therefore, creating a semantic search engine that can be used easily by individuals with varying familiarity with mathlib4 is very important. In this paper, we present a semantic search engine333Our search engine is expected to launch within a month, available as a cloud service. for mathlib4 that accepts informal queries and finds the relevant theorems. We also establish a benchmark for assessing the performance of various search engines for mathlib4.
keywords:
Semantic Search, Machine Learning, Benchmarkscategory:
[…]
@@
Lean [2, 3] is an interactive theorem prover built on dependent type theory, designed to verify mathematical proofs written in a formal language and thereby enhancing their rigor. It has a vibrant and supportive community, with its popularity growing among mathematicians. A prime example of the Lean community’s collaborative spirit is mathlib4, the mathematical library for Lean 4. This library, regularly updated by contributors from around the globe, acts as a basis for the formalization of new mathematical theories. This eliminates the need to repeatedly formalize established results, as users can simply check if mathlib4 contains the necessary theorems referenced in informal proofs. However, locating these theorems is often challenging due to the limitations of officially provided search tools, which struggle to find relevant theorems with informal queries.
@@
There are primarily two methods to search for theorems in mathlib4: using the mathlib4 documentation444https://leanprover-community.github.io/mathlib4_docs/ and searching the GitHub repository of mathlib4555https://github.com/leanprover-community/mathlib4. The mathlib4 documentation allows users to search for theorems by their formal names. However, this feature can be challenging for beginners to utilize effectively, as they may not be familiar with the naming conventions. For example, Cauchy’s Mean Value Theorem is named as exists_ratio_deriv_eq_ratio_slope in mathlib4, meaning a direct search for "Cauchy’s Mean Value Theorem" yields no results. An alternative method involves searching within the GitHub repository of mathlib4, which allows for keyword-based searches across the source files, including formal statements, proofs, and documentation strings. This method can locate Cauchy’s Mean Value Theorem as it is mentioned in the documentation string of exists_ratio_deriv_eq_ratio_slope. However, this approach faces two issues: many theorems in mathlib4 lack documentation strings, and semantically similar user queries that don’t exactly match the theorems or documentation strings may lead to unsuccessful searches.
@@
Consequently, neither method adequately supports finding theorems based on informal queries, leading to beginners spending significant time on this task. According to a survey among students working on formalizing convex analysis at Peking University, they spend half of their time searching for theorems. Moreover, conversations on Zulip666https://leanprover.zulipchat.com/#narrow/stream/219941-Machine-Learning-for-Theorem-Proving/topic/Semantic.20Search.20for.20Mathematics have highlighted the need for creating a semantic search engine for mathlib4. Consequently, the development of such a search engine for mathlib4 is highly desirable to enhance the efficiency of theorem retrieval.
@@ 2 Related Work
Text retrieval is the task of finding relevant information within a corpus based on user queries. Early methods, such as BM25 [26, 25], used sparse vector representations for queries and documents, assessing relevance by comparing these vectors with certain weighting techniques [27]. While effective in measuring lexical similarity, these approaches fall short in capturing the semantic similarity between queries and documents.
To address this limitation, deep learning techniques [6, 33, 17, 14] have been introduced. Utilizing deep neural networks, these techniques encode queries and documents into dense vectors, thereby assessing relevance based on semantic similarity. Further developments have been made in neural architectures and training paradigms. There are two main architectures: the cross-encoder [23, 22] and the bi-encoder [9, 24, 21, 20]. Cross-encoders take the concatenation of query and document as input and produce the final relevance of this query-document pair, while bi-encoders map the query and document into vectors separately, determining relevance through similarity between the two vectors. Training paradigms have also evolved, with the Inverse Close Task (ICT) [12] initially proposed for dense retriever pre-training. Subsequently, other pre-training tasks have been developed, including Body First Selection and Wiki Link Prediction, both introduced in [1]. Recent studies [18, 30, 28, 32] have explored large-scale unsupervised pre-training using contrastive loss, followed by fine-tuning on smaller, labeled datasets. Pre-training with extensive text pairs allows the language model to grasp textual semantics and the fine-tuning stage further enhances its performance across various retrieval tasks. However, the authors of [31] argue that a two-stage training approach might not be necessary, demonstrating that directly fine-tuning a decoder-only model on both synthetic and labeled data can yield competitive results.
@@
Mathematical Information Retrieval (MIR) differs from text retrieval in that it involves queries and documents that contain mathematical formulas. These formulas are highly structured, which distinguishes them from plain text. To effectively capture the semantics of math formulas, several representations are employed. The Symbol Layout Tree (SLT) [35] preserves the original layout of formulas, while the Operator Tree (OPT) [5] represents mathematical symbols as nodes, with edges denoting the relationships between operators and operands. Classical MIR methods [4, 5, 11, 36, 8, 19, 34] rely on structure search, identifying matching substructures across various features. Among these, the Approach0 structural search method [41, 38] has shown to be particularly effective. It indexes formulas through leaf-root paths in the OPT and utilizes subexpression matching to assess formula similarity. With advancements in deep learning, dense retrievers have been integrated into MIR, often combined with structural searches [8, 39, 40, 37]. A notable example is the Approach0 hybrid search [39], which combines Approach0 structure search with a bi-encoder dense retriever, ColBERT [10]. This combination not only facilitates effective formula matching but also enhances understanding of context.
The preceding literature focuses on extracting mathematical content from a corpus composed of natural language texts and formulas formatted in markup languages. This contrasts with our objective of conducting searches within mathlib4, a formal mathematical library. The work most closely related to ours is Moogle777https://www.moogle.ai/, a semantic search engine for mathlib4. However, its technical details have not been released. We will compare the performance of Moogle and our search engine in Section 5.
@@ 3 Methodology
In this section, we will describe the implementation of our semantic search engine for mathlib4. This engine is designed to accept a user query in natural language and return a list of relevant theorems in mathlib4. Our approach involves converting formal theorems from mathlib4 into their informal counterparts, as illustrated in Figure 1. These informal-formal theorem pairs are then vectorized and stored in a vector database, a step that can be executed offline. When a user query is submitted, we augment the query to better grasp its context, vectorize the enriched query, and locate the corresponding theorems in the embedding space. In the following subsections, we will elaborate on the informalization of mathlib4, the design of the search engine, and the method of query augmentation.
@@ 3.1 Informalizing Mathlib4
Notably, the authors of [7] also employed an LLM for informalizing mathlib4 statements, relying solely on the formal statements. We argue that providing the LLM with additional context, such as related definitions and documentation strings, enhances its ability to accurately interpret and convert formal theorems into their informal counterparts.
@@ 3.2 Semantic Search Engine for Mathlib4
After obtaining the informal corpus, we employ dense embedding models, which excel at capturing semantic information, to encode the informal-formal theorem pairs. Recent advancements in text embedding models have introduced the practice of integrating specific task instructions into either queries or documents, enhancing the model’s adaptability to diverse tasks and boosting performance in zero-shot settings [28, 13, 31]. Consequently, for the purpose of theorem retrieval within mathlib4, we enrich our corpus documents with specific instructions. We adopt the following instruction template:
@@
These task-specific instructions significantly influence the performance of embedding models. We will investigate the effect of varying task instructions on the overall retrieval performance in Section 5.3.1.
Embedding the entire corpus, although a time-consuming process, is performed offline and does not require repetition for each use of the engine. We utilize Chroma DB to store the embeddings. Upon receiving a query, our system vectorizes it and retrieves theorems based on their cosine distance to the query in the embedding space. Chroma DB employs the Hierarchical Navigable Small World (HNSW) algorithm [15], an efficient approximate nearest neighbor search method, ensuring rapid retrieval from the corpus.
@@ 3.3 Query Augmentation
To enhance query clarity and achieve a more accurate embedding, our system incorporates a query augmentation step. As illustrated in Figure 4, this process involves prompting an LLM101010We use gpt-4-0125-preview for query augmentation. to transform a brief, vague query into a detailed statement that includes both informal and formal statements, ensuring mathematical equivalence with the original query. We guide the LLM with specific principles for query augmentation, emphasizing the importance of precision, the use of LaTeX for mathematical expressions, and the clarification of ambiguous inputs. Additionally, we provide examples of query augmentations to improve the LLM’s comprehension of the task. Although Lean 4 code generation by LLMs may occasionally introduce inaccuracies due to the limited presence of mathlib4 in their training data, this approach effectively enriches the query with additional contextual information.
@@ 4 Mathlib4 Semantic Search Benchmark
To rigorously assess and compare the efficacy of various retrieval methods, we have established the Mathlib4 semantic search benchmark. This benchmark encompasses a curated set of queries, relevance labels of each each Mathlib4 theorem for these queries, and a collection of performance metrics. In this section, we will explain this benchmark in detail. Section 4.1 describes the composition of the query set. In Section 4.2, we explain the relevance criteria for our benchmark along with the whole labeling procedure. Section 4.3 lists the performance metrics used in our benchmark.
@@ 4.1 Composition of the Query set
The query set of our benchmark has 50 distinct queries, spanning various mathematical disciplines including calculus, abstract algebra, linear algebra, number theory, algebraic number theory, set theory, and mathematical logic. This selection aims to cover a broad spectrum of topics and complexities. We argue that this is a common size for mathematical information retrieval database, as evidenced by the ARQMath1, ARQMath2, and ARQMath3 databases from the MIR field, which contain 77, 71, and 78 queries, respectively, for their answer retrieval tasks [16].
To optimize labeling efforts, we organize queries with identical search intents into 18 distinct, non-overlapping groups, each containing at least two distinct queries. This approach assumes that all queries within a group have the same relevance score for any document and significantly reduces the need for repetitive labeling, as each document is evaluated just once per query group. Furthermore, to provide a more detailed assessment and mitigate the impact of duplicate document labels, we consider four prevalent forms of mathematical queries: natural language descriptions, LaTeX formulas, theorem names, and Lean 4 term descriptions. In each query group, we strive to include as many different description forms as possible. Table 1 provides statistics and examples of the query set. We note that not all query groups have all four different representation forms, as it is often the case that a theorem may not have an official name or its representation in a LaTeX formula might be redundant.
@@
Natural Description |
18 |
If there exist injective maps of sets from to and from to , then there exists a bijective map between and . |
If implies , then not implies not . |
LaTeX Formula |
15 |
If there exist injective, injective, then there exists bijective. |
Theorem Name |
7 |
Schroeder Bernstein Theorem |
Modus Tollens |
Lean 4 Term |
10 |
{f : A → B} {g : B → A} (hf : Injective f) (hg : Injective g) : h, Bijective h | (p → q) → (¬q → ¬p) |
@@ 4.2 Relevance Judgments
During the process of collecting relevance labels for query-document pairs, we adopt an approach similar to ARQMath[16]. Initially, we establish a carefully crafted relevance assessment criteria, elaborated in Table 2. While doing labeling, assessors are instructed to evaluate whether a given search result facilitates their mathematical formalization workflow. Instead of simply considering similarity in topic, presented form, formula structure or mathematical induction relationship, the relevance we consider here are deeply engaged with Lean 4 experts’ need.
@@
Exact Match |
2 |
1 |
Exact match to the query or being a stronger statement |
Relevant |
1 |
0.3 |
Useful in locating where the corresponding statement should be |
Irrelevant |
0 |
0 |
Not expected to be useful in mathematics formalization workflow |
For each set of queries grouped by identical search intentions, assessors are presented with the top 50 theorems list retrieved by an intermediate version of our search engine. In addition to evaluating the provided theorems, assessors are tasked with identifying and adding any relevant theorems that may have been omitted from the initial list by inspecting the files where "Exact Match" items in the list are located, with particular attention to those with "Exact Match" rating.
Given the structured organization of Mathlib4, where related theorems are often located within the same file, it is reasonable to assume that any items not in the list are irrelevant to this query group. We also note that all queries are assured to have at least one exact match in Mathlib4. These two assumptions support the performance metrics we use in the benchmark, as described in the following subsection.

### v2

@@ Abstract
The interactive theorem prover Lean enables the verification of formal mathematical proofs and is backed by an expanding community. Central to this ecosystem is its mathematical library, mathlib4, which lays the groundwork for the formalization of an expanding range of mathematical theories. However, searching for theorems in mathlib4 can be challenging. To successfully search in mathlib4, users often need to be familiar with its naming conventions or documentation strings. Therefore, creating a semantic search engine that can be used easily by individuals with varying familiarity with mathlib4 is very important. In this paper, we present a semantic search engine111The search engine is available at https://leansearch.net. for mathlib4 that accepts informal queries and finds the relevant theorems. We also establish a benchmark for assessing the performance of various search engines for mathlib4.222The data and the source code are available at https://github.com/reaslab/LeanSearch.
A Semantic Search Engine for Mathlib4
Guoxiong Gao1††thanks: These authors contributed equally., Haocheng Ju211footnotemark: 1, Jiedong Jiang2, Zihan Qin1, Bin Dong,3,4††thanks: Corresponding author 1School of Mathematical Sciences, Peking University, Beijing, China 2Beijing International Center for Mathematical Research, Peking University, Beijing, China 3Beijing International Center for Mathematical Research and the New Cornerstone Science Laboratory, Peking University, Beijing, China 4Center for Machine Learning Research, Peking University, Beijing, China samggx@stu.pku.edu.cn hcju@pku.edu.cn {emailboxofjjd,15261375660}@163.com dongbin@math.pku.edu.cn
1 Introduction
@@
Lean (de Moura et al., 2015; de Moura and Ullrich, 2021) is an interactive theorem prover built on dependent type theory, designed to verify mathematical proofs written in a formal language and thereby enhancing their rigor. It has a vibrant and supportive community, with its popularity growing among mathematicians. A prime example of the Lean community’s collaborative spirit is mathlib4, the mathematical library for Lean 4. This library, regularly updated by contributors from around the globe, acts as a basis for the formalization of new mathematical theories. This eliminates the need to repeatedly formalize established results, as users can simply check if mathlib4 contains the necessary theorems referenced in informal proofs. However, locating these theorems is often challenging due to the limitations of officially provided search tools, which struggle to find relevant theorems with informal queries.
@@
There are primarily two methods to search for theorems in mathlib4: using the mathlib4 documentation333https://leanprover-community.github.io/mathlib4_docs/ and searching the GitHub repository of mathlib4444https://github.com/leanprover-community/mathlib4. The mathlib4 documentation allows users to search for theorems by their formal names. However, this feature can be challenging for beginners to utilize effectively, as they may not be familiar with the naming conventions. For example, Cauchy’s Mean Value Theorem is named as exists_ratio_deriv_eq_ratio_slope in mathlib4, meaning a direct search for "Cauchy’s Mean Value Theorem" yields no results. An alternative method involves searching within the GitHub repository of mathlib4, which allows for keyword-based searches across the source files, including formal statements, proofs, and documentation strings. This method can locate Cauchy’s Mean Value Theorem as it is mentioned in the documentation string of exists_ratio_deriv_eq_ratio_slope. However, this approach faces two issues: many theorems in mathlib4 lack documentation strings, and semantically similar user queries that don’t exactly match the theorems or documentation strings may lead to unsuccessful searches.
@@
Consequently, neither method adequately supports finding theorems based on informal queries, leading to beginners spending significant time on this task. Discussions on Zulip555https://leanprover.zulipchat.com/#narrow/stream/219941-Machine-Learning-for-Theorem-Proving/topic/Semantic.20Search.20for.20Mathematics have highlighted the need for creating a semantic search engine for mathlib4. Therefore, developing such a search engine for mathlib4 is highly desirable to improve the efficiency of theorem retrieval.
@@ 2 Related Work
Text retrieval is the task of finding relevant information within a corpus based on user queries. Early methods, such as BM25 (Robertson et al., 1995; Robertson and Zaragoza, 2009), used sparse vector representations for queries and documents, assessing relevance by comparing these vectors with certain weighting techniques (Salton and Buckley, 1988). While effective in measuring lexical similarity, these approaches fall short in capturing the semantic similarity between queries and documents.
To address this limitation, deep learning techniques (Huang et al., 2013; Xiong et al., 2017; McDonald et al., 2018; Liu et al., 2018) have been introduced. Utilizing deep neural networks, these techniques encode queries and documents into dense vectors, thereby assessing relevance based on semantic similarity. Further developments have been made in neural architectures and training paradigms. There are two main architectures: the cross-encoder (Qiao et al., 2019; Nogueira et al., 2020) and the bi-encoder (Karpukhin et al., 2020; Qu et al., 2021; Ni et al., 2022b, a). Cross-encoders take the concatenation of query and document as input and produce the final relevance of this query-document pair, while bi-encoders map the query and document into vectors separately, determining relevance through similarity between the two vectors. Training paradigms have also evolved, with the Inverse Close Task (ICT) (Lee et al., 2019) initially proposed for dense retriever pre-training. Subsequently, other pre-training tasks have been developed, including Body First Selection and Wiki Link Prediction, both introduced in (Chang et al., 2020). Recent studies (Neelakantan et al., 2022; Wang et al., 2022; Su et al., 2023; Xiao et al., 2023) have explored large-scale unsupervised pre-training using contrastive loss, followed by fine-tuning on smaller, labeled datasets. Pre-training with extensive text pairs allows the language model to grasp textual semantics and the fine-tuning stage further enhances its performance across various retrieval tasks. However, the authors of (Wang et al., 2024) argue that a two-stage training approach might not be necessary, demonstrating that directly fine-tuning a decoder-only model on both synthetic and labeled data can yield competitive results.
@@
Mathematical Information Retrieval (MIR) differs from text retrieval in that it involves queries and documents that contain mathematical formulas. These formulas are highly structured, which distinguishes them from plain text. To effectively capture the semantics of math formulas, several representations are employed. The Symbol Layout Tree (SLT) (Zanibbi and Blostein, 2012) preserves the original layout of formulas, while the Operator Tree (OPT) (Gao et al., 2016) represents mathematical symbols as nodes, with edges denoting the relationships between operators and operands. Classical MIR methods (Fraser et al., 2018; Gao et al., 2016; Kristianto et al., 2016; Zanibbi et al., 2016; Kane et al., 2022; Ng et al., 2021, 2020) rely on structure search, identifying matching substructures across various features. Among these, the Approach0 structural search method (Zhong and Zanibbi, 2019; Zhong et al., 2020) has shown to be particularly effective. It indexes formulas through leaf-root paths in the OPT and utilizes subexpression matching to assess formula similarity. With advancements in deep learning, dense retrievers have been integrated into MIR, often combined with structural searches (Kane et al., 2022; Zhong et al., 2022a, b, 2023). A notable example is the Approach0 hybrid search (Zhong et al., 2022a), which combines Approach0 structure search with a bi-encoder dense retriever, ColBERT (Khattab and Zaharia, 2020). This combination not only facilitates effective formula matching but also enhances understanding of context.
The preceding literature focuses on extracting mathematical content from a corpus composed of natural language texts and formulas formatted in markup languages. This contrasts with our objective of conducting searches within mathlib4, a formal mathematical library. The work most closely related to ours is Moogle666https://www.moogle.ai/, a semantic search engine for mathlib4. However, its technical details have not been released. We will compare the performance of Moogle and our search engine in Section 5.
@@ 3 Methodology
In this section, we will describe the implementation of our semantic search engine for mathlib4777We use mathlib4 with commit db04a978b67b2200691f0bd968f334c83261b66a.. This engine is designed to accept a user query in natural language and return a list of relevant theorems in mathlib4. Our approach involves converting formal theorems from mathlib4 into their informal counterparts, as illustrated in Figure 1. These informal-formal theorem pairs are then vectorized and stored in a vector database, a step that can be executed offline. When a user query is submitted, we augment the query to better grasp its context, vectorize the enriched query, and locate the corresponding theorems in the embedding space. In the following subsections, we will elaborate on the informalization of mathlib4, the design of the search engine, and the method of query augmentation.
@@ 3.1 Informalizing Mathlib4
Notably, the authors of (Jiang et al., 2023) also employed an LLM for informalizing mathlib4 statements, relying solely on the formal statements. We argue that providing the LLM with additional context, such as related definitions and documentation strings, enhances its ability to accurately interpret and convert formal theorems into their informal counterparts. An example comparing the informalized mathlib4 statement generated by (Jiang et al., 2023) to the one produced by our method is provided in Appendix A.
@@ 3.2 Semantic Search Engine for Mathlib4
After obtaining the informal corpus, we employ dense embedding models, which excel at capturing semantic information, to encode the informal-formal theorem pairs. Recent advancements in text embedding models have introduced the practice of integrating specific task instructions into either queries or documents, enhancing the model’s adaptability to diverse tasks and boosting performance in zero-shot settings (Su et al., 2023; Li and Li, 2023; Wang et al., 2024). Consequently, for the purpose of theorem retrieval within mathlib4, we enrich our corpus documents with specific instructions. We adopt the following instruction template:
@@
These task-specific instructions significantly influence the performance of embedding models. We will investigate the effect of varying task instructions on the overall retrieval performance in Appendix D.
Embedding the entire corpus, although a time-consuming process, is performed offline and does not require repetition for each use of the engine. We utilize Chroma DB to store the embeddings. Upon receiving a query, our system vectorizes it and retrieves theorems based on their cosine distance to the query in the embedding space. Chroma DB employs the Hierarchical Navigable Small World (HNSW) algorithm (Malkov and Yashunin, 2020), an efficient approximate nearest neighbor search method, ensuring rapid retrieval from the corpus.
@@ 3.3 Query Augmentation
To enhance query clarity and achieve a more accurate embedding, our system incorporates a query augmentation step. This process involves prompting an LLM101010We use gpt-4-0125-preview for query augmentation. to transform a brief, vague query into a detailed statement that includes both informal and formal statements, ensuring mathematical equivalence with the original query. We guide the LLM with specific principles for query augmentation, emphasizing the importance of precision, the use of LaTeX for mathematical expressions, and the clarification of ambiguous inputs. Additionally, we provide examples of query augmentations to improve the LLM’s comprehension of the task. A detailed prompt for query augmentation is provided in Appendix B. Although Lean 4 code generation by LLMs may occasionally introduce inaccuracies due to the limited presence of mathlib4 in their training data, this approach effectively enriches the query with additional contextual information.
@@ 4 Mathlib4 Semantic Search Benchmark
To rigorously assess and compare the efficacy of various retrieval methods, we have established the Mathlib4 Semantic Search Benchmark. This benchmark encompasses a curated set of queries, relevance labels of each mathlib4 theorem for these queries, and a collection of performance metrics. In this section, we provide a detailed explanation of the benchmark. Section 4.1 describes the composition of the query set. In Section 4.2, we explain the relevance criteria for our benchmark and the labeling procedure. Section 4.3 lists the performance metrics used in our benchmark.
@@ 4.1 Composition of the Query Set
The query set of our benchmark has 50 distinct queries, spanning various mathematical disciplines including calculus, abstract algebra, linear algebra, number theory, algebraic number theory, set theory, and mathematical logic. This selection aims to cover a broad spectrum of topics and complexities. We argue that this is a typical size for the query sets in mathematical information retrieval datasets, as demonstrated by the ARQMath1, ARQMath2, and ARQMath3 databases from the MIR field, which contain 77, 71, and 78 queries, respectively, for their answer retrieval tasks (Mansouri et al., 2022).
To optimize labeling efforts, we organize queries with identical search intents into 18 distinct, non-overlapping groups, each containing at least two distinct queries. This approach assumes that all queries within a group have the same relevance score for any document, significantly reducing the need for repetitive labeling, as each document is evaluated only once per query group. Furthermore, to provide a more detailed assessment and mitigate the impact of duplicate document labels, we consider four prevalent forms of mathematical queries: natural language descriptions, LaTeX formulas, theorem names, and Lean 4 term descriptions. In each query group, we strive to include as many different description forms as possible. Table 1 provides statistics and examples of the query set. However, not all query groups contain all four representation forms, as some theorems may not have an official name or their representation in a LaTeX formula might be redundant.
@@
| Natural Description | 18 | If there exist injective maps of sets from to and from to , then there exists a bijective map between and . | If implies , then not implies not . |
| LaTeX Formula | 15 | If there exist injective, injective, then there exists bijective. | |
| Theorem Name | 7 | Schroeder Bernstein Theorem | Modus Tollens |
| Lean 4 Term | 10 | {f : A → B} {g : B → A} (hf : Injective f) (hg : Injective g) : h, Bijective h | (p → q) → (¬q → ¬p) |
@@ 4.2 Relevance Judgments
During the process of collecting relevance labels for query-document pairs, we adopt an approach similar to ARQMath(Mansouri et al., 2022). Initially, we establish a carefully crafted relevance assessment criteria, elaborated in Table 2. While performing the labeling, assessors are instructed to evaluate whether a given search result facilitates their mathematical formalization workflow. Instead of merely considering similarity in topic, presented form, formula structure, or mathematical induction relationship, the relevance we focus on here is deeply aligned with the needs of Lean 4 experts.
@@
| Exact Match | 2 | 1 | Exact match to the query or being a stronger statement |
| Relevant | 1 | 0.3 | Useful in locating where the corresponding statement should be |
| Irrelevant | 0 | 0 | Not expected to be useful in mathematics formalization workflow |
For each set of queries grouped by identical search intentions, assessors are presented with the top 50 theorems retrieved by an intermediate version of our search engine. In addition to evaluating the provided theorems, assessors are tasked with identifying and adding any relevant theorems that may have been omitted from the initial list by inspecting the files where "Exact Match" items in the list are located, with particular attention to those with "Exact Match" rating.
Given the structured organization of mathlib4, where related theorems are often located within the same file, it is reasonable to assume that any items not in the list are irrelevant to this query group. Additionally, all queries are guaranteed to have at least one exact match in mathlib4. These assumptions support the performance metrics we use in the benchmark, as described in the following subsection.